OpenStudy (karatechopper):

State the interval over which f(x)=-6x^2+12x-4 is decreasing. A) (-infinity, -1) B) (-1, infinity) C) (1, infinity) D) (-infinity, 1)

OpenStudy (freckles):

This question looks incomplete.

OpenStudy (freckles):

State the interval over which the function what?

OpenStudy (karatechopper):

Ah I see, fixed it!

OpenStudy (karatechopper):

Refresh

OpenStudy (freckles):

Ok.

OpenStudy (freckles):

Do you know calculus?

OpenStudy (freckles):

We can also do this with algebra.

OpenStudy (freckles):

The trick is finding the vertex.

OpenStudy (karatechopper):

I know pre-calculus. Algebra is preferred.

OpenStudy (karatechopper):

Yess the vertex, I'm having issues with that. I tried to complete the square.

OpenStudy (anonymous):

Graph this using this site https://www.desmos.com/calculator It should be (1, +infinity)

OpenStudy (karatechopper):

No, I don't like depending on sites to graph things out for me. I want to know how to do it algebraically.

OpenStudy (karatechopper):

But I appreciate it.

OpenStudy (anonymous):

well then use factoring, and then find the midpoint of the two zeros and plug the midpoint into the function and find the max.

OpenStudy (freckles):

\[f(x)=ax^2+bx+c \\ f(x)=ax^2+\frac{a}{a}bx+c \\ f(x)=a(x^2+\frac{1}{a}bx)+c \\ f(x)=a(x^2+\frac{b}{a}x)+c \\f(x)=a(x^2+\frac{b}{a}x+( \frac{b}{2a} )^2)+c-a ( \frac{b}{2a})^2 \\ f(x)=a(x+\frac{b}{2a})^2+c-a \frac{b^2}{4a^2} \\ f(x)=a(x+\frac{b}{2a})^2+c-\frac{b^2}{4a} \\ f(x)=a(x+\frac{b}{2a})^2+\frac{4ac-b^2}{4a}\]

OpenStudy (anonymous):

We know it will be max because the degree is even, and leading coeifficient is negative, so the parabola will look like this. 

OpenStudy (karatechopper):

Ok ok hold on.

OpenStudy (freckles):

Therefore the vertex is \[(\frac{-b}{2a},\frac{4ac-b^2}{4a})\] But we really only care about the x-coordinate of the vertex which is x=-b/(2a)

OpenStudy (karatechopper):

Before we move on. Can I show you what I did?

OpenStudy (freckles):

Ok I will hold on. lol.

OpenStudy (anonymous):

yes

OpenStudy (karatechopper):

Ok, would be it correct to complete the square?

OpenStudy (freckles):

Yes! You can follow that scheme I laid out above.

OpenStudy (freckles):

That involves completing the square for any quadratic

OpenStudy (karatechopper):

Freckles, yours is a bit different than mine. Let me show you how I learned it.

OpenStudy (freckles):

Alright.

OpenStudy (karatechopper):

-6x^2+12x-4 -6(x^2-2x+____)=4+_____ (b/2)^2 (-2/2)^2=-1^2=1 -6(x^2-2x+1)=4+6 -6(x^2-2x+1)=10 -6(x-1)^2=10 -6(x-1)^2-10 Oh LOL I did it wrong on paper then while typing it out I got it I think. -10 is vertex?!

OpenStudy (karatechopper):

oh wait no

OpenStudy (karatechopper):

(1,-10) is vertex

OpenStudy (karatechopper):

Right?

OpenStudy (freckles):

Well I guess by you mean you did it different you mean put an equal sign and you didn't add and take away on the same side but you added on both sides . \[-6x^2+12x-4=f(x) \\-6(x^2-2x)-4=f(x) \\ -6(x^2-2x+1)-4=f(x)-6 \\ -6(x-1)^2-4=f(x)-6 \\ -6(x-1)^2-4+6=f(x)\]

OpenStudy (freckles):

so basically I'm saying you subtracted on one side but added on the other side

OpenStudy (freckles):

you added -6 on left hand side and added +6 on the other side

OpenStudy (karatechopper):

Ah I see it would be -6(x-1)^2-2

OpenStudy (karatechopper):

So the vertex becomes (1,-2)

OpenStudy (freckles):

Well the vertex is (1,-4+6)

OpenStudy (freckles):

So I'm saying -6(x-1)^2+2=f(x)

OpenStudy (freckles):

and therefore the vertex is (1,2)

OpenStudy (freckles):

But anyways for your problem we only care about the x-coordinate of the vertex.

OpenStudy (freckles):

What other good info you can determine which interval it is decreasing on even increasing on is to determine if the parabola is open up or open down

OpenStudy (karatechopper):

OHH I SEEEEEEEEE WHY ITS 2 POSTIVE LOL

OpenStudy (karatechopper):

Right ok.

OpenStudy (karatechopper):

its open down

OpenStudy (freckles):

Since our a is negative, then our parabola is open down which mean the function is increasing on (-inf,1) and decreasing on (1,inf)

OpenStudy (karatechopper):

So it would be decreasing from 1 to the right? so 1 to infinity?

OpenStudy (karatechopper):

(1,infinity)?

OpenStudy (karatechopper):

High five!

OpenStudy (freckles):

High five!

OpenStudy (karatechopper):

Thank you very much for helping me! :)xxxxx

OpenStudy (freckles):

Np.

OpenStudy (karatechopper):

Have a great day wherever!

OpenStudy (anonymous):

You know a easier method would be just to get the derivative of the function, and make it equal zero to get the value of x for the max. and just plug it into the function and you get the Max point. I dont know why i didnt think of that before lol.